A solution reaches a concentration of 0.014 molar N2 in water. What is the applied pressure in torr of the gas? (Use 6.25*10-4 M/atm as the Henry's law constant kH for N2)
Answer #1

Calculate the applied pressure as follows:

M = kH*P

Then, P = 0.014/(6.25*10^-4)

P = 22.4 atm

Using, 1 atm = 760 torr

Then, P = 22.4*760 = 17024 torr

Therefore, the required answer is 17024 torr.

Please give good rating.